Dylan is 9 years older than Parry. Together their ages sum to 65 years. How old is Parry?

Answer by Saleye(10)

Let 'd' & 'p' be the ages of of Dylan & Parry respectively:
We've: =>Dylan is 9 years older than Parry.
=>
d=p+9 ---(eq. 1)
Also sum of their age:
=> d+p=65---(eq. 2)

=>Subtitute 'd' from eq. 1 in to two:
=> (p+9)+p=65
solve for p:
p=28 years <= the age of Parry.
Age of Dylan
d=p+9=37